OpenStudy (anonymous):

Dividing Radical ∛(a^3+b^3 )/∛(a+b)

OpenStudy (anonymous):

Radicals are just expressions for powers, so you can use power laws to turn your orginal expression into:\[\left( \frac{a^3+b^3}{a+b} \right)^{1/3}\]

OpenStudy (anonymous):

You then use polynomial or synthetic division to divide that expression.

OpenStudy (anonymous):

The fraction is\[a^2-ab+b^2\]so your expression will become,\[\sqrt[3]{a^2-ab+b^2}\]

OpenStudy (anonymous):

How you division

OpenStudy (anonymous):

Hmm, it's a bit difficult to explain on this site. I'll see if I can find a link with a clip.

OpenStudy (anonymous):

It's not a hard thing to do - you'll pick it up quickly.

OpenStudy (anonymous):

can you tye wgat number I need division

OpenStudy (anonymous):

Sorry, I'm sure what you want?

OpenStudy (anonymous):

You still want me to do the division?

OpenStudy (anonymous):

division like this: a+b:〖 a〗^3+ b^3

OpenStudy (anonymous):

you just write eqution and I know how to division

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I just want to know what number division by what number

OpenStudy (anonymous):

a+b | a^3 + b^3

OpenStudy (anonymous):

The denominator appears on the LHS, the numerator on the RHS.

OpenStudy (anonymous):

let my try,thank

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

Can any of u help me how to draw the graph

OpenStudy (anonymous):

The graph of what?

OpenStudy (anonymous):

thank, I can solve now

OpenStudy (anonymous):

No probs. Become a fan :)